4.

In a study, the following equation for predicting a child’s metabolic rate (in 100 kcal/24h) (Y) given his/her weight (in kilograms) (X) is found: Y = 0.4X + 0.89. The RMS for the prediction is 1.144 (in 100 kcal/24h). What is the probability that a child that weighs 15 kilograms has metabolic rate exceeding 8?

Question 4 options:

A)

0.834

B)

Can’t tell without the data

C)

0.144

D)

0.166

E)

0.970

5.

Suppose a study predicting calories (Y) based on sugar content (X) in breakfast cereals produced the following information. Calculate the regression equation for this dataset:
Ux=7.1 x=4.5
Jy=107.2 Oү=19.6     R=.564

Question 5 options:

A)

slope: 3.315, intercept: 29.477

B)

slope: 6.037, intercept: 9.474

C)

slope: 1.386, intercept: 13.241

D)

slope: 2.457, intercept: 89.755

E)

slope: 5.981, intercept: 30.152

6.

Suppose you want to predict the final exam score of a student given that their number of hours studying for this class is at 80th percentile. Assume both final exam scores and number of hours studying for this class are normally distributed. If the correlation coefficient is 0.85, what would you expect for their final exam score based on the regression effect?

Question 6 options:

A)

They will score above the 80th percentile for their final exam.

B)

They will score below the 80th percentile for their final exam.

C)

They will score exactly at the 80th percentile for their final exam.

D)

They will score exactly at the 85th percentile for their final exam.

E)

We do not have enough information to answer this question.
Answer #1

4)

probability that a child that weighs 15 kilograms has metabolic rate exceeding 8 =P(X>8)=P(Z>(8-0.4*15-0.89)/1.144)

=P(Z>0.97) =0.166

option D is correct

5)

x̅ = 7.1
y̅ = 107.2
sx= 4.5
sy = 19.6
r= 0.564
slope b1=r*sy/sx= 2.457
intercept bo=y̅ +r(sy/sx)*(x̅)= 89.759

option D is correct

6)option B is correct

They will score below the 80th percentile for their final exam.